# Group inverses of matrices of simple symmetric directed graphs

### v1

@@ Abstract.
A new class of simple symmetric digraphs called $\mathbb{D}$ is defined and studied here. Any digraph in $\mathbb{D}$ has the property that each non-pendant is adjacent to at least one pendant vertex. A graph theoretical description for the entries of the group inverse of a real square matrix $A$ with any digraph belonging to this class is given. We classify all the real square matrix $A$ such that the digraphs associated with $A$ and $A^{\#}$ both are in $\mathbb{D}$, that is the digraph related to $A$ is either a corona or a star digraph.
@@
aaa
@@ 1. Introduction
Firstly, let us recall the definition of the group inverse of a matrix, the object of primary interest here. For a real $n\times n$ matrix $A$, the group inverse, if it exists, is the matrix $X$ that satisfies the equations $AXA=A,XAX=X$ and $AX=XA$. Such an $X$ is always unique and denoted by $A^{\#}$. Any matrix $X$ that satisfies the first equation is called an inner inverse of $A$ while satisfying the second equation will be called an outer inverse. It is well-known that the group inverse of $A$ exists if and only if $rank(A)=rank(A^{2})$. For a non-singular matrix $A$, $A^{-1}=A^{\#}$. Let us recall that for a real rectangular matrix $A$, the Moore-Penrose inverse of $A$, is the unique matrix $A^{\dagger}$ that satisfies the equations $AA^{\dagger}A=A,~{}A^{\dagger}AA^{\dagger}=A^{\dagger},~{}(AA^{\dagger})^{T}=AA^{\dagger}$ and $(A^{\dagger}A)^{T}=A^{\dagger}A$. We refer the reader to [7] for more details on these notions of generalized inverses and Moore-Penrose inverses.
Now, recall some graph-theoretic notation from [15]. Let $A=(a_{ij})$ be an $n\times n$ matrix with real entries. The digraph $D(A)=(V,E)$ corresponding to $A$ is the directed graph whose vertex set is $V=\{1,2,\ldots,n\}$ and whose edge set $E$ is described by the requirement that $(i,j)\in E$ iff $a_{ij}\neq 0$. For $m\geq 1$, a sequence $(i_{1},i_{2},\ldots,i_{m},i_{m+1})$ of $m+1$ distinct vertices with edges $(i_{1},i_{2}),(i_{2},i_{3}),\ldots,(i_{m},i_{m+1})$ in $E$ is called a path of length $m$ from $i_{1}$ to $i_{m+1}$ in $D(A)$. For $m\geq 2$, a sequence $(i_{1},i_{2},\ldots,i_{m},i_{1})$ with distinct $m$ vertices $i_{1},i_{2},\ldots,i_{m},$ where $(i_{1},i_{2}),(i_{2},i_{3}),\ldots,(i_{m},i_{1})\in E,$ is called an $m$-$cycle$ (a cycle of length $m$) in $D(A)$.
A real square matrix $A=(a_{ij})$ is called combinatorially symmetric if $a_{ij}\neq 0$ iff $a_{ji}\neq 0$. Trivially, any symmetric matrix is combinatorially symmetric. Parter was the first author to use the concept of combinatorially symmetric matrices to study important problems significantly [24]. In 1974, Maybee first introduced the notion of a combinatorially symmetric matrix in [21]. Digraph $D(A)$ corresponding to a matrix $A$ is called a simple symmetric digraph if it has no loops and there is an edge in each direction between two distinct vertices i.e. the corresponding matrix $A$ is combinatorially symmetric matrix with zero diagonal. Recall that a directed graph that has a path from each vertex to every other vertex is called strongly connected. Digraph $D(A)$ corresponding to a matrix $A$ is called a tree if it is strongly connected, and all of its cycles have length 2. Clearly, a tree digraph is a simple symmetric digraph.
Suppose $A$ is a real square matrix such that $D(A)$ is a simple symmetric digraph. A cycle $(i,j,i)$ is said to be incident to $i$ as well as $j$ in $D(A)$. A vertex $i$ is called a pendant vertex if it is incident to only one 2-cycle and non-pendant vertex if it is incident to more than one 2-cycle in $D(A)$. A 2-cycle $(i,j,i)$ will be called a pendant cycle if at least one vertex $i$ or $j$ is pendant in $D(A)$, while a 2-cycle which is not pendant will be called a non-pendant cycle. A pair of vertices $i,j$ is said to be adjacent to each other if there is a 2-cycle $(i,j,i)$ in $D(A)$. For $r$ even, a set of $\frac{r}{2}$ disjoint 2-cycles in $D(A)$ given by $\{(i_{1},i_{2},i_{1}),(i_{3},i_{4},i_{3}),\ldots,(i_{r-1},i_{r},i_{r-1})\},$ is called a matching and the product $a_{i_{1},i_{2}}a_{i_{2},i_{1}}a_{i_{3},i_{4}}a_{i_{4},i_{3}}\ldots a_{i_{r-1},i_{r}}a_{i_{r},i_{r-1}}$ is called a matching product. If this set of 2-cycles has a maximum cardinality then the matching is referred to as a maximum matching and the matching product is then called a maximum matching product. The sum of all maximum matching products in $D(A)$ is denoted by $\Delta_{A}$. A matching is said to be a perfect matching if it covers all the vertices of $D(A)$. $\mathbb{M}$ and $\mathbb{M}(i)$ denote the set of all maximum matchings in $D(A)$ and the set of all maximum matchings in which vertex $i$ is matched, respectively. For a cycle $(i,j,i)$ in $D(A)$, the product $a_{ij}a_{ji}$ is called the cycle product. A sequence of $m$ 2-cycles $((i_{1},i_{2},i_{1}),(i_{2},i_{3},i_{2}),\ldots,(i_{m},i_{m+1},i_{m}))$ with $m+1$ distinct vertices $i_{1},i_{2},\ldots,i_{m+1}$ in $D(A)$ is called a cycle chain from $i_{1}$ to $i_{m+1}$ of length $m$ and denoted by $C_{m}(i_{1},i_{m+1})$. A cycle chain $C_{m}(i_{1},i_{m+1})$ is said to be an alternating cycle chain with respect to a maximum matching $M$ if cycles of $C_{m}(i_{1},i_{m+1})$ alternatively belong to $M$ and $M^{c}$, with the condition that both the first and the last cycle of $C_{m}(i_{1},i_{m+1})$ belong to $M$. For an alternating cycle chain $C_{m}(i_{1},i_{m+1})$, the product $a_{i_{1},i_{2}}a_{i_{2},i_{3}}a_{i_{3},i_{4}}\ldots a_{i_{m-1},i_{m}}a_{i_{m},i_{m+1}}$ is said to be the path product along alternating cycle chain $C_{m}(i_{1},i_{m+1})$, denoted by $P_{m}(i_{1},i_{m+1})$.
@@
In combinatorial matrix theory, an interesting problem is giving the inverse or group inverse of a matrix using its graph structure. Here are some articles that explain how to determine the inverse [1], [4], [5], [13], [20], [25] and the group inverse [8], [12], [15], [16], [17], [23], [27]. In chemistry, the group inverse of adjacency matrices is used to determine a separation gap of a molecule by using graph energies [26]. An important parameter in quantum chemistry is the smallest positive eigenvalue. There appear to be fewer methods for computing the smallest positive eigenvalue of a matrix in contrast to a large number of techniques for computing the largest positive eigenvalue (even approximately). For a square matrix $A$, it is well known that the set of all the non-zero eigenvalues of $A^{\#}$ is the set of all reciprocals of the non-zero eigenvalues of $A$. Thus, just as is the case for nonsingular matrices, to obtain estimates for the smallest positive eigenvalue of $A$, one may consider estimates for the largest positive eigenvalue of $A^{\#}$. Also, the group inverses of matrices have applications in algebraic connectivity [19], Markov chains [11] and resistance distance [10].
@@
We recall some well-known results on the role of generalized inverses in graph theory. In [15, Theorem 2.2], authors computed the group inverse of the bipartite matrix of the form $A=\begin{pmatrix}0&B\\ C&0\end{pmatrix}$ and presented a graphical description for the entries of $A^{\#}$ for path digraph $D(A)$. In the same article, they proposed a conjecture ([15, Conjecture 5.1]) on the entries of $A^{\#}$ when $D(A)$ is a tree digraph and this conjecture is solved recently in [23, Theorem 2.11] for a special class of trees. A block-wise description for $A^{\#}$ was given in [16, Theorem 2.1] when $B,C$ of the form $[X\ U],[Y\ V]^{T}$ respectively, with $X,Y$ non-singular and $rank(UV)=1$. Also, a graphical description for the entries of $A^{\#}$ such that $D(A)$ is a broom tree was derived in [16]. In [27], the authors gave a graphical description for entries of the adjacency matrix of arbitrary weighted trees in terms of maximum matchings and alternating paths. Group inverses matrices associated with cycle graphs were investigated in [14, 17, 26, 28].
Organization of the paper: In section 2, we define a new class $\mathbb{D}$ of simple symmetric directed graphs. We recall a block-wise description for the group inverse of a matrix $A$ with $D(A)\in\mathbb{D}$, then provide a graph theoretical description for the entries of $A^{\#}$ in Theorem 2.8. In section 3, we characterize all the real square matrix $A$ such that $D(A)$ and $D(A^{\#})$ both are in $\mathbb{D}$ in Theorem 3.5.
2. Graphical Description of $A^{\#}$ for a Matrix with Simple Symmetric Digraph
In [15], authors presented a graphical description for the entries of group inverse of a matrix $A$ with path digraph $D(A)$ and proposed a conjecture for tree digraph. Recently, in [23], we gave a formula for the entries of the group inverse of a matrix with a special class of directed trees and proved that the conjecture proposed in [15] holds for this special class of directed trees. In this section, we define a new class of simple symmetric directed graphs $\mathbb{D}$ which is an extension of the class considered in [23, Definition 1.2] and showed that the formula obtained in [23, Theorem 1.4] is extendable for the class $\mathbb{D}$ (Theorem 2.8).
Let us now introduce a new class of directed graphs.
@@ 2. Graphical Description of $A^{\#}$ for a Matrix with Simple Symmetric Digraph
Let $\mathbb{D}$ denote the set of all simple symmetric directed graphs $D$ such that each non-pendant vertex of $D$ is adjacent to at least one pendant vertex of $D$.
An important point to note is that the underlying graphs (of this class $\mathbb{D}$) are a special case of cluster networks derived by taking arbitrary graphs as bases and stars as satellites, [3, 30]. Cluster networks are highly relevant in applications in Chemistry since all composite molecules consisting of some amalgamation over a central submolecule can be understood as generalized cluster networks. For instance, they can be used to understand some issues in metal-metal interaction in some molecules, since a cluster network structure can be easily found. In [30], the Kirchhoff index formulae for composite graphs known as join, corona and cluster of two graphs, are presented, in terms of the pieces. The Kirchhoff index formulae and the effective resistances of generalized composite networks, such as generalized cluster or corona network are obtained, in terms of the pieces, in [3]. We refer the reader to [2, 6] for more details on the kirchhoff index of networks.
@@
It is clear that the digraph $D_{1}\in\mathbb{D},$ (Fig. 1) while $D_{2}\notin\mathbb{D}$ (Fig. 2). The non-pendant vertex $4$ (in $D_{2}$) is not adjacent to any pendant vertex.
@@
In another way, the property of a digraph $D$ in $\mathbb{D}$ can be written as follows: If $P\subset V(D)$ is the set of pendant vertices, then $\delta(P)=V(D)\backslash P$ or in other words, $P$ is a dense subset of $V(D)$. Sometimes, this property is referred to as the set $P$ dominates $V(D)$. Remember that for a set $F\subset V(D)$ we have $\overline{F}=F\cup\delta(F)$, where $\delta(F)=\{x\in V(D)\backslash F:x\mbox{ is adjacent to }y\mbox{ for some }y\in F\}$. Suppose a directed graph $D$ is in $\mathbb{D}$.
Recall that a directed graph that has a path from each vertex to every other vertex is called strongly connected. A square matrix $A$ is said to be irreducible if $D(A)$ is strongly connected. Let $N(i)$ denote the set of all vertices adjacent to vertex $i$. Suppose a directed graph $D$ is in $\mathbb{D}$. Then, each component of $D$ also belongs to $\mathbb{D}$. Henceforth, we will consider a strongly connected graph in $\mathbb{D}$. Now, we obtain some special properties for the digraphs in $\mathbb{D}$ in the following few results. In general, they are not valid for any simple symmetric digraph.
Proposition 2.3.
Let $D\in\mathbb{D}$ be a strongly connected graph. Then, no non-pendant cycle can belong to a maximum matching of D.
Proof.
In [23, Proposition 2.2], we prove the same result for a tree digraph $D\in\mathbb{D}$. For any digraph $D\in\mathbb{D}$, we can obtain this result by giving the same proof. ∎
Remark 2.4.
Let $D\in\mathbb{D}$ have $k$ non-pendant vertices. Then, the number of 2-cycles in a maximum matching is always $k$. Note that, every non-pendant vertex is matched in any maximum matching of $D$, and both the endpoints of a length three alternating cycle chain are pendant vertices and a length one alternating cycle chain is nothing but a pendant cycle.
Proposition 2.5.
Let $D\in\mathbb{D}$ be a strongly connected graph. We then have the following:
$(i)$
The length of any alternating cycle chain is at most three.
$(ii)$ Any alternating cycle chain between any two vertices is unique.
Proof.
$(ii)$ Let $C$ be an alternating cycle chain from vertex $i$ to $j$ in $D$. Then, by first part, the length of $C$ is at most three. If the length is one then $C$ is simply a pendant cycle. So, it is unique. Now, let us consider the case when the length of $C$ is three. Since pendant cycles are the only maximum matching cycles (corresponding to any maximum matching), the initial and the terminal vertices of $C$ should be pendant vertices. So, $i$ and $j$ are pendant vertices. Let $\tilde{C}$ be another alternating cycle chain from $i$ to $j$. Since $i$ and $j$ are pendant vertices, the initial and the terminal 2-cycles of $C$ and $\tilde{C}$ must be the same. Here, $\tilde{C}$ is an alternating cycle chain, so its length should be odd. $C$ and $\tilde{C}$ are two different alternating cycle chains with the condition that they have the same initial and terminal 2-cycles. This means that the length of the cycle chain $\tilde{C}$ should be at least five, a contradiction to the fact that the length of any alternating cycle chain in $D$ is at most three (as we proved in part (i)). Therefore, $C$ and $\tilde{C}$ coincide. ∎
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Suppose $D(A)$ has $k$ non-pendant vertices. For $i=1,2,\ldots,k$, let $x_{i}$ and $y_{i}$ be (column) vectors of length $r_{i}\in\mathbb{N}$ such that every coordinate is nonzero $(\sum_{i=1}^{k}r_{i}=n-k)$. We shall refer to such vectors as strictly nonzero vectors. Set
Then $A$ can be written permutationally similar to a matrix of the form
$\begin{pmatrix}E&F\\ G&0\\ \end{pmatrix},$ (2.1)
where $E=(e_{ij})\in\mathbb{R}^{k\times k}$. Set $\alpha_{i}=x_{i}^{T}y_{i}$ for all $i\in\{1,2,\ldots,k\}$. Let $W=(W_{ij})$ be the $k\times k$ block matrix, where $W_{ij}=(\frac{e_{ij}}{\alpha_{i}\alpha_{j}})y_{i}x_{j}^{T}$, an $r_{i}\times r_{j}$ matrix. Then, by [22, Theorem 2.5], the group inverse of matrices of the form 2.1 has the $2\times 2$ block form
$\begin{pmatrix}0&~{}Y\\ Z&-W\\ \end{pmatrix},$ (2.2)
For a square matrix $A$, $A^{\#}$ may not always exist. For instance, if $A=\begin{pmatrix}0&1\\ 0&0\\ \end{pmatrix},$ then $A^{\#}$ does not exist. In particular, it is easy to show that the group inverse does not exist for any nilpotent matrix. The following proposition provides a necessary and sufficient condition for the existence of the matrix $A^{\#}$ when $D(A)\in\mathbb{D}$, that is the sum of all maximum matchings in $D(A)$ is nonzero.
Proposition 2.6.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is strongly connected. Then $A^{\#}$ exists if and only if $\Delta_{A}\neq 0$.
Proof.
$A$ has the form 2.1 since $D(A)\in\mathbb{D}$. Now, set $\alpha_{i}=x_{i}^{T}y_{i}$ for all $i\in\{1,2,\ldots,k\}$. For each $i$, the non-pendant vertex $i$ is adjacent to $r_{i}$ pendant vertices and $\alpha_{i}$ is the sum of cycle products of $r_{i}$ pendant cycles incident $i$. Note that
By [29, Theorem 3.1], $A^{\#}$ exists if and only if $rank(F)=rank(FG)$. Here, $rank(F)=k$ which implies $A^{\#}$ exists if and only if $\alpha_{i}\neq 0$ for all $i\in\{1,2,\ldots,k\}$. By Proposition 2.3, all the maximum matching cycles are pendant cycles and a maximum matching of $D(A)$ has a set of $k$ pendant cycles incident to $k$ non-pendant vertices (the total number of maximum matching is $\prod_{i=1}^{k}r_{i}$). So, $\Delta_{A}=\prod_{i=1}^{k}\alpha_{i}$ and it follows that $A^{\#}$ exists if and only if $\Delta_{A}\neq 0$. ∎
@@
Let us define more terminology for a strongly connected graph $D(A)\in\mathbb{D}$, where $A$ is a real square matrix. For arbitrary vertices $i$ and $j$ in $D(A)$ denote $\mathbb{M}(i,j)$ to be the set of all maximum matchings $M$ in $D(A)$ such that $C_{m}(i,j)$ is an alternating cycle chain with respect to $M$. Clearly, $\mathbb{M}(i,j)=\mathbb{M}(j,i)$. A necessary condition for the set $\mathbb{M}(i,j)$ to be non-empty is that the length of the path from $i$ to $j$ be odd. If $(i,j,i)$ is a 2-cycle of some maximum matching, then $\mathbb{M}(i,j)$ is non-empty. Two distinct vertices $i$ and $j$ will be called maximally matchable if $\mathbb{M}(i,j)\neq\phi$ and by Proposition 2.5, for a maximally matchable vertices $i,j$, $C_{m}(i,j)$ is unique.
@@
In this terminology, a formula for the entries of the group inverse of a matrix $A$ with tree digraph $D(A)\in\mathbb{D}$, proved in [23], is recalled next.
@@
[23, Theorem 1.4] Let $A$ be an $n\times n$ real matrix with a tree graph $D(A)\in\mathbb{D}$ and assume that $\Delta_{A}\neq 0$. Let $A^{\#}=(\alpha_{ij})$ and let $\mu_{ij}$ be defined as above. Then, $\alpha_{ij}=\frac{\mu_{ij}}{\Delta_{A}}$.
Here is the main result of this section:
@@
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is a strongly connected graph and $\Delta_{A}\neq 0$. Let $A^{\#}=(\alpha_{ij})$ and let $\mu_{ij}$ be defined in 2.3. Then, $\alpha_{ij}=\frac{\mu_{ij}}{\Delta_{A}}$.

### v3

# Matrices with simple symmetric digraphs and their group inverses

@@ Abstract.
A new class of simple symmetric digraphs called $\mathcal{D}$ is defined and studied here. Any digraph in $\mathcal{D}$ has the property that each non-pendant vertex is adjacent to at least one pendant vertex. A graph theoretical description for the entries of the group inverse of a real square matrix with any digraph belonging to this class is given. We classify all the real square matrices $A$ such that the digraphs associated with $A$ and $A^{\#}$ both are in $\mathcal{D}$, that is, the digraph related to $A$ is either a corona or a star digraph.
@@
Group inverse; Simple symmetric digraph; Maximum matching; Alternating cycle chain.
@@ 1. Introduction
Firstly, let us recall the definition of the group inverse of a matrix, the object of primary interest here. For a real $n\times n$ matrix $A$, the group inverse, if it exists, is the matrix $X$ that satisfies the equations $AXA=A,XAX=X$ and $AX=XA$. Such an $X$ is always unique and denoted by $A^{\#}$. It is well-known that the group inverse of $A$ exists if and only if $rank(A)=rank(A^{2})$. For a non-singular matrix $A$, $A^{-1}=A^{\#}$. Let us recall that for a real rectangular matrix $A$, the Moore-Penrose inverse of $A$, is the unique matrix $A^{\dagger}$ that satisfies the equations $AA^{\dagger}A=A,~{}A^{\dagger}AA^{\dagger}=A^{\dagger},~{}(AA^{\dagger})^{T}=AA^{\dagger}$ and $(A^{\dagger}A)^{T}=A^{\dagger}A$. We refer the reader to [7] for more details on these notions of generalized inverses and Moore-Penrose inverses.
@@
In combinatorial matrix theory, an interesting problem is giving the inverse or group inverse of a matrix using its graph structure. Here are some articles that explain how to determine the inverse [1], [4], [5], [12], [19], [24] and the group inverse [8], [11], [14], [15], [16], [22], [26]. In chemistry, the group inverse of adjacency matrices is used to determine the separation gap of a molecule using graph energies [25]. In addition, the group inverses of the matrices have applications in algebraic connectivity [18], Markov chains [10] and resistance distance [9].
@@
We recall some well-known results on the role of generalized inverses in graph theory. In [14, Theorem 2.2], authors computed the group inverse of the bipartite matrix of the form $A=\begin{pmatrix}0&B\\ C&0\end{pmatrix}$ and presented a graphical description for the entries of $A^{\#}$ for path digraph $D(A)$. In the same article, they proposed a conjecture ([14, Conjecture 5.1]) on the entries of $A^{\#}$ when $D(A)$ is a tree digraph and recently, we solved this conjecture for a special class of trees in [22, Theorem 2.11]. A block-wise description for $A^{\#}$ was given in [15, Theorem 2.1] when $B,C$ of the form $[X\ U],[Y\ V]^{T}$ respectively, with $X,Y$ non-singular and $rank(UV)=1$. Also, a graphical description for the entries of $A^{\#}$ such that $D(A)$ is a broom tree was derived in [15]. In [26], the authors gave a graphical description for entries of the adjacency matrix of arbitrary weighted trees in terms of maximum matchings and alternating paths. Group inverses matrices associated with cycle graphs were investigated in [13, 16, 25, 27].
Now, we recall some graph-theoretic notations coined from [14]. Let $A=(a_{ij})\in\mathbb{R}^{n\times n}$. Then, the digraph corresponding to $A$, denoted by $D(A)=(V,E)$, is the directed graph whose vertex set is $V=\{1,2,\ldots,n\}$ and the edge set $E$ is defined as follows: $(i,j)\in E$ iff $a_{ij}\neq 0$. A sequence $(i_{1},i_{2},\ldots,i_{m},i_{m+1})$ of $m+1$ distinct vertices with edges $(i_{1},i_{2}),(i_{2},i_{3}),\ldots,(i_{m},i_{m+1})$ is called a path of length $m$ from vertex $i_{1}$ to vertex $i_{m+1}$ in $D(A)$. When the first and last vertex of the above sequence are the same, i.e. $i_{1}=i_{m+1}$, it is called a $m$-$cycle$ (a cycle of length $m$) in $D(A)$.
A digraph $D$ is called a simple symmetric digraph if it has no loops and there is an edge in each direction between two distinct vertices. Recall that a digraph that has a path from each vertex to every other vertex is called strongly connected. A digraph is said to be tree if it is a strongly connected digraph, and all of its cycles have length 2. Clearly, a tree digraph is a simple symmetric digraph. Suppose $A$ is a real square matrix such that $D(A)$ is a simple symmetric digraph. A 2-cycle $(i,j,i)$ is said to be incident to $i$ as well as $j$ in $D(A)$, and the vertex $i$ is called a pendant vertex if it is incident to only one 2-cycle. Otherwise, we will call $i$ a non-pendant vertex. On the other hand, a 2-cycle $(i,j,i)$ will be called a pendant cycle if at least one vertex $i$ or $j$ is pendant in $D(A)$, while a 2-cycle which is not pendant will be called a non-pendant cycle. A pair of vertices $i,j$ is said to be adjacent to each other if there is a 2-cycle $(i,j,i)$ in $D(A)$.
Now, extending the notations in [22], for a simple symmetric digraph $D(A)$, recall that for an even $r$, a set of $\frac{r}{2}$ disjoint 2-cycles in $D(A)$ given by $\{(i_{1},i_{2},i_{1}),(i_{3},i_{4},i_{3}),\ldots,(i_{r-1},i_{r},i_{r-1})\},$ is called a matching and the product $a_{i_{1},i_{2}}a_{i_{2},i_{1}}a_{i_{3},i_{4}}a_{i_{4},i_{3}}\ldots a_{i_{r-1},i_{r}}a_{i_{r},i_{r-1}}$ is called a matching product. If this set of 2-cycles has a maximum cardinality then the matching is referred to as a maximum matching and the matching product is called a maximum matching product. The sum of all maximum matching products in $D(A)$ is denoted by $\Delta_{A}$. A matching is said to be a perfect matching if it covers all the vertices of $D(A)$. Let $\mathbb{M}$ and $\mathbb{M}(i)$ denote the set of all maximum matchings in $D(A)$ and the set of all maximum matchings in which vertex $i$ is matched, respectively. For a cycle $(i,j,i)$ in $D(A)$, the product $a_{ij}a_{ji}$ is called the cycle product. A sequence of $m$ 2-cycles $((i_{1},i_{2},i_{1}),(i_{2},i_{3},i_{2}),\ldots,(i_{m},i_{m+1},i_{m}))$ with $m+1$ distinct vertices $i_{1},i_{2},\ldots,i_{m+1}$ in $D(A)$ is called a cycle chain from $i_{1}$ to $i_{m+1}$ of length $m$ and denoted by $C_{m}(i_{1},i_{m+1})$. A cycle chain $C_{m}(i_{1},i_{m+1})$ is said to be an alternating cycle chain with respect to a maximum matching $M$ if cycles of $C_{m}(i_{1},i_{m+1})$ alternatively belong to $M$ and $M^{c}$, with the condition that both the first and the last cycle of $C_{m}(i_{1},i_{m+1})$ belong to $M$. For an alternating cycle chain $C_{m}(i_{1},i_{m+1})$, the product $a_{i_{1},i_{2}}a_{i_{2},i_{3}}a_{i_{3},i_{4}}\ldots a_{i_{m-1},i_{m}}a_{i_{m},i_{m+1}}$ is said to be the path product along alternating cycle chain $C_{m}(i_{1},i_{m+1})$, denoted by $P_{m}(i_{1},i_{m+1})$.
Let us introduce a new class of simple symmetric digraph.
@@
Let $\mathcal{D}$ denote the set of all simple symmetric digraphs $D$ such that each non-pendant vertex of $D$ is adjacent to at least one pendant vertex of $D$.
@@
It is clear that the digraph $D_{1}\in\mathcal{D},$ (Fig. 1) while $D_{2}\notin\mathcal{D}$ (Fig. 2). The non-pendant vertex $4$ (in $D_{2}$) is not adjacent to any pendant vertex.
@@
In another way, the property of a digraph $D$ in $\mathcal{D}$ can be written as follows: If $P\subset V(D)$ is the set of pendant vertices, then $\delta(P)=V(D)\backslash P$ or in other words, $P$ is a dense subset of $V(D)$. Sometimes, this property is referred to as the set $P$ dominates $V(D)$. Remember that for a set $F\subset V(D)$ we have $\overline{F}=F\cup\delta(F)$, where $\delta(F)=\{x\in V(D)\backslash F:x\mbox{ is adjacent to }y\mbox{ for some }y\in F\}$.
@@
Let us define more terminology for a strongly connected digraph $D(A)\in\mathcal{D}$, where $A$ is a real square matrix. In particular, when $D(A)$ is a tree digraph, these terminologies are defined in [22]. Now, we are extending those terminologies for a simple symmetric digraph $D(A)\in\mathcal{D}$. For arbitrary vertices $i$ and $j$ in $D(A)$ denote $\mathbb{M}(i,j)$ to be the set of all maximum matchings $M$ in $D(A)$ such that $C_{m}(i,j)$ is an alternating cycle chain with respect to $M$. Clearly, $\mathbb{M}(i,j)=\mathbb{M}(j,i)$. A necessary condition for the set $\mathbb{M}(i,j)$ to be non-empty is that the length of the path from $i$ to $j$ be odd. If $(i,j,i)$ is a 2-cycle of some maximum matching, then $\mathbb{M}(i,j)$ is non-empty. Two distinct vertices $i$ and $j$ will be called maximally matchable if $\mathbb{M}(i,j)\neq\phi$ and by Proposition 2.3, for a maximally matchable vertices $i,j$, $C_{m}(i,j)$ is unique.
@@
In this terminology, a formula for the entries of the group inverse of a matrix $A$ with tree digraph $D(A)\in\mathcal{D}$, proved in [22], is recalled next.
@@
[22, Theorem 1.4] Let $A$ be an $n\times n$ real matrix with a tree digraph $D(A)\in\mathcal{D}$ and assume that $\Delta_{A}\neq 0$. Let $A^{\#}=(\alpha_{ij})$ and let $\mu_{ij}$ be defined as above. Then, $\alpha_{ij}=\frac{\mu_{ij}}{\Delta_{A}}$.
In [14], authors presented a graphical description for the entries of group inverse of a matrix $A$ with path digraph $D(A)$ and proposed a conjecture for tree digraph. Recently, in Theorem 1.3, we gave a formula for the entries of the group inverse of a matrix with a special class of tree digraphs and proved that the conjecture proposed in [14] holds for this special class of tree digraphs. Our first main result shows that [22, Theorem 1.4] is extendable for the larger class $\mathcal{D}$.
@@
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathcal{D}$. Assume that $D(A)$ is a strongly connected digraph and $\Delta_{A}\neq 0$. Let $A^{\#}=(\alpha_{ij})$ and let $\mu_{ij}$ be defined in 1.1. Then, $\alpha_{ij}=\frac{\mu_{ij}}{\Delta_{A}}$.
Example 3.3 shows that the group inverse of a real square matrix $A$ with a simple symmetric digraph may not be a matrix with a simple symmetric digraph, whereas $D(A^{\#})$ is a simple symmetric digraph if $D(A)\in\mathcal{D}$. Now, a natural question arises: When $D(A)$ and $D(A^{\#})$ both are in $\mathcal{D}$ for a real square matrix $A$? We give the answer in our final main result.
Theorem 1.5.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathcal{D}$. Assume that $D(A)$ is strongly connected and $\Delta_{A}\neq 0$. Then $D(A^{\#})\in\mathcal{D}$ if and only if $D(A)$ is either a corona digraph or a star tree digraph.
Organization of the paper: The remaining sections are devoted to proving our main results above. In section 2, we give a few properties of digraphs in $\mathcal{D}$ and recall a block-wise description for the group inverse of a matrix $A$ with $D(A)\in\mathcal{D}$, and prove Theorem 1.4 by giving three Lemmas. In section 3, we provide a necessary and sufficient condition for the $ij$-th entry of the matrix $A^{\#}$ to be non-zero for a real square matrix $A$ with $D(A)\in\mathcal{D}$ and prove Theorem 1.5.
2. Graphical description for the group inverses of matrices with simple symmetric digraphs
Recall that a real square matrix $A=(a_{ij})$ is called combinatorially symmetric if $a_{ij}\neq 0$ iff $a_{ji}\neq 0$. Trivially, any symmetric matrix is combinatorially symmetric. Parter was the first author to use the concept of combinatorially symmetric matrices [23]. In 1974, Maybee first introduced the notion of a combinatorially symmetric matrix in [20]. Note that for a simple symmetric digraph $D(A)$, the corresponding matrix $A$ is a combinatorially symmetric matrix with zero diagonal.
An important point to note is that the underlying graphs (of this class $\mathcal{D}$) are a special case of cluster networks derived by taking arbitrary graphs as bases and stars as satellites, [3, 29]. Cluster networks are highly relevant in applications in Chemistry since all composite molecules consisting of some amalgamation over a central submolecule can be understood as generalized cluster networks. For instance, they can be used to understand some issues in metal-metal interaction in some molecules, since a cluster network structure can be easily found. In [29], the Kirchhoff index formulae for composite graphs known as join, corona and cluster of two graphs, are presented, in terms of the pieces. The Kirchhoff index formulae and the effective resistances of generalized composite networks, such as generalized cluster or corona network are obtained, in terms of the pieces, in [3]. We refer the reader to [2, 6] for more details on the kirchhoff index of networks.
A real square matrix $A$ is said to be irreducible if $D(A)$ is strongly connected. Let $N(i)$ denote the set of all vertices adjacent to vertex $i$. Suppose a digraph $D$ is in $\mathcal{D}$. Then, each component of $D$ also belongs to $\mathcal{D}$. Henceforth, we will consider a strongly connected digraph in $\mathcal{D}$. Now, we obtain some special properties for the digraphs in $\mathcal{D}$ in the following few results. In general, they are not valid for any simple symmetric digraph.
Proposition 2.1.
Let $D\in\mathcal{D}$ be a strongly connected digraph. Then, no non-pendant cycle can belong to a maximum matching of D.
Proof.
In [22, Proposition 2.2], we prove the same result for a tree digraph $D\in\mathcal{D}$. For any digraph $D\in\mathcal{D}$, we can obtain this result by giving the same proof. ∎